OpenStudy (anonymous):

plz help i will give a medal and fan What is the slope-intercept form of the function that contains the point (–1, 5) and has a slope of –4?

rishavraj (rishavraj):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m\]

OpenStudy (anonymous):

oh ok but about what -4 there only one number

rishavraj (rishavraj):

it would be \[\frac{ y - 5 }{ x + 1 } = -4\]

OpenStudy (anonymous):

y=mx+b m is the slope and b is the y intercept